# Bayesian Invariance Modeling of Multi-Environment Data

### v1

@@ C.2 Uncertainty quantification in limited environments
Effect of estimation bias.
In the above examples, we observe that the posterior does not assign equal mass to multiple invariant features, despite that a uniform prior is used. This result is an artifact of estimation bias arising from finite data, which will disappear when the true models are available.
@@ C.3.1 Simulation details
We describe the process to create a random set of joint distributions $\{p_{e}(x,y)\}_{e=1}^{E}$ given $p$ and $E$. The reason to differentiate between various $p$ is to control the scale of parameter values to prevent the sampled $x,y$ values from exploding when $p$ increases.
@@
      $actprob\sim\textrm{Uniform}\left(\{0.6,0.7,0.8,0.9\}\right)$
@@
      For $e=1$: (i) coefficient lower bound $lb=1$; (ii) coefficient upper bound $ub=2.1$, (iii) noise level lower bound $\sigma_{\min}=0.1$; (iv) noise level upper bound $\sigma^{2}_{\max}=0.2$.
@@
      Separately for $e=2,\cdots,E$: for the intervened variable (i) mean of the absolute value of the intercept $m_{e}\sim\textrm{Uniform}\left([0,1]\right)$; (ii) coefficient lower bound $lb_{e}=(m_{e}+0.01)*2$; (iii) coefficient upper bound $ub_{e}=(m_{e}+0.5)*2$; (iv) noise level lower bound $\sigma_{e,\min}\sim\textrm{Uniform}\left([0.1,0.2]\right)$; (v) noise level upper bound $\sigma_{e,\max}=\sigma_{e,\min}+\textrm{Uniform}\left([0.1,0.5]\right)$; (vi)the probability of changing the coefficient $changeprob_{e}\sim\textrm{Uniform}\left([0.1,0.3]\right)$; and (vii) the fraction of features to be intervened on $intfrac_{e}\sim\textrm{Uniform}\left([0.5,1]\right)$
@@
      $actprob\sim\textrm{Uniform}\left(\{0.1,0.15,0.2,0.25\}\right)$
@@
      Separately for $e=2,\cdots,E$: for the intervened variable (i) mean of the absolute value of the intercept $m_{e}\sim\textrm{Uniform}\left([0,0.4]\right)$; (ii) coefficient lower bound $lb_{e}=m_{e}+0.01$; (iii) coefficient upper bound $ub_{e}=(m_{e}+0.01)*1.5$; other hyperparameters are chosen in the same way as $p=10$ case
@@ C.3.2 Method implementation details
    VI-BIP: The prior is uniform over $p=450$$\{z\in\{0,1\}^{p}:\|z\|_{0}\leq p_{\max}\}$ where $p_{\max}=10$ in low-dimensional case ($p=10$), and $p_{\max}=450$ in high dimensions $(p=450)$.

### v2

@@ C.3.1 Simulation details
We follow the general data generative process described in the preamble of Supplementary Section C with the following configuration (The reason to differentiate between various $p$ is to control the scale of parameter values to prevent the sampled $x,y$ values from exploding when $p$ increases.):
@@
      $p_{\text{act}}\sim\textrm{Uniform}\left(\{0.6,0.7,0.8,0.9\}\right)$
@@
      For $e=1$: (i) coefficient lower bound $lb=1$; (ii) coefficient upper bound $ub=2.1$, (iii) noise level lower bound $\sigma_{\min}=0.1$; (iv) noise level upper bound $\sigma_{\max}=0.2$.
@@
      Separately for $e=2,\cdots,E$: for the intervened variable (i) mean of the absolute value of the intercept $m_{e}\sim\textrm{Uniform}\left([0,1]\right)$; (ii) coefficient lower bound $lb_{e}=(m_{e}+0.01)*2$; (iii) coefficient upper bound $ub_{e}=(m_{e}+0.5)*2$; (iv) variance multiplicative factor lower bound $\lambda_{e,\min}\sim\textrm{Uniform}\left([0.1,0.2]\right)$; (v) variance multiplicative factor relative range $\lambda_{e,\textrm{diff}}\sim\textrm{Uniform}\left([0.1,0.5]\right)$; (vi)the probability of changing the coefficient $p_{\text{change}}\sim\textrm{Uniform}\left([0.1,0.3]\right)$; and (vii) the fraction of features to be intervened on $\rho^{\text{int}}_{e}\sim\textrm{Uniform}\left([0.5,1]\right)$
@@
      $p_{\text{act}}\sim\textrm{Uniform}\left(\{0.1,0.15,0.2,0.25\}\right)$
@@
      Separately for $e=2,\cdots,E$: for the intervened variable (i) mean of the absolute value of the intercept $m_{e}\sim\textrm{Uniform}\left([0,0.4]\right)$; (ii) coefficient lower bound $lb_{e}=m_{e}+0.01$; (iii) coefficient upper bound $ub_{e}=(m_{e}+0.01)*1.5$; other simulation hyperparameters are chosen in the same way as $p=10$ case
@@ C.3.2 Method implementation details
    VI-BIP: The prior is uniform over $\{z\in\{0,1\}^{p}:\|z\|_{0}\leq p_{\max}\}$. In low dimensions $(p=10)$, $p_{\max}=10$ leading to an uninformative prior; In high dimensions ($p=450$), $p_{\max}=20$ encouraging sparsity.